OpenStudy (anonymous):

if cos (theta) = 7/25 and in quadrant one how would I find cos2 (theta)

OpenStudy (phi):

are you asking about cos squared \( \cos^2(\theta) \) or cos of 2 theta \( \cos(2 \theta) \) ?